OpenStudy (anonymous):

Verify the identity. cos4u=cos^22u−sin^22u

OpenStudy (anonymous):

This is: \[ \cos(4u) = \cos^2(2u)-\sin^2(2u) \]Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you use: \[ \cos(x+y) = \cos(x)\cos(y) - \sin(x)\sin(y) \]??

OpenStudy (anonymous):

I think so, thats what Im confused about. cos(2u+2u)=cos(2u)sin(2u)-sin(2u)cos(2u) then what?

OpenStudy (anonymous):

What bout using: \[ \cos^2(x) = \frac{1+\cos(2x)}{2} \]and: \[ \sin^2(x) = \frac{1-\cos(2x)}{2} \]

OpenStudy (anonymous):

NOT cos(2u)sin(2u)-sin(2u)cos(2u) IT'S cos(2u)cos(2u)-sin(2u)sin(2u)

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

If you're asking that question, then you don't understand. If you don't understand, don't write something down as if you do.

OpenStudy (raden):

use the identity : cos(2x) = cos^2 x - sin^2 x now, cos4u = cos2(2u) = .....

OpenStudy (anonymous):

Im writing it as a question because Im asking

OpenStudy (anonymous):

There a many identities you can use. You just have to do the algebra right.

OpenStudy (anonymous):

which means I dont understand get it?

OpenStudy (anonymous):

Once you get it like this: \[ \cos(2u)\cos(2u)-\sin(2u)\sin(2u) \] Then it becomes this: \[ \cos^2(2u)-\sin^2(2u) \]

OpenStudy (anonymous):

cos2(2u) = cos(2u)cos(2u)-sin(2u)sin(2u) so thats it then?

OpenStudy (raden):

cos * cos = cos^2 , @Anna4061

OpenStudy (raden):

same for sin

OpenStudy (anonymous):

No, I'm saying: \[ \cos(2u) \cos(2 u) = \cos^2(2u) \]and likewise with sin

OpenStudy (anonymous):

I get that, but what do I do after? thats all?

OpenStudy (raden):

i think u have done it, and wio is right cos2(2u) = cos(2u)cos(2u)-sin(2u)sin(2u) cos(4u) = cos^2 (2u) - sin^2 (2u)

OpenStudy (anonymous):

ok thanks

OpenStudy (raden):

yw